# Circle Squaring with Pieces of Small Boundary and Low Borel Complexity

### v3

@@ 2.5. Some Auxiliary Structures in $G_{d}$
At certain points during the construction, we are forced to make arbitrary “tie-breaking” choices in a consistent manner. In doing so, it is useful to endow ${\mathbbm{T}}^{k}$ with a partial order in which each component of $G_{d}$ is totally ordered. We will use the lexicographic ordering where, for distinct $\boldsymbol{u},\boldsymbol{v}\in\mathbbm{T}^{k}$, we write $\boldsymbol{u}\prec_{\operatorname{lex}}\boldsymbol{v}$ to mean $\boldsymbol{u}]_{G_{d}}=[\boldsymbol{v}]_{G_{d}}$ and the first non-zero entry of $\vec{n}$ is positive, where $\vec{n}$ is the (unique) element of ${\mathbbm{Z}}^{d}$ with
@@
Note that the families of subsets of ${\mathbbm{T}}^{k}$ that are Jordan measurable, or have boundary of upper Minkowski dimension at most some given constant, are both examples of $a$-invariant algebras.
@@
For any $r\geqslant 0$ there exists $\delta>0$ such that every set $X\subseteq\mathbbm{T}^{k}$ whose projection onto the first coordinate has diameter at most $\delta$ is $r$-discrete.
@@
Since $X=C_{1}^{\prime}\cup\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt\cup C_{m}^{\prime}$, it is enough to show that each $C_{i}^{\prime}$ is an r(i1)r(i-1)-local function of $C_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,C_{i}$. We use induction on $i=1,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,m$. This is true for $i=1$ since the set $C_{1}^{\prime}=C_{1}$ is a 0-local function of $C_{1}$. For each $2\leqslant i\leqslant m$, the set $C_{i}^{\prime}$ is a Boolean combination of $C_{i}$ and the sets $C_{1}^{\prime},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,C_{i-1}^{\prime}$ translated by vectors of the form $\sum_{i=1}^{d}n_{i}\boldsymbol{x}_{i}$ where $\|\vec{n}\|_{\infty}\leqslant r$. So, by induction, $C_{i}^{\prime}$ is an r(i1)r(i-1)-local function $C_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,C_{i}$, as desired.
@@ 2.7. Equidecompositions from Integer-Valued Flows
Start by taking $S_{\boldsymbol{t}_{1}}$ to be a strip of width $\gamma$ such that the intersection $A_{\boldsymbol{t}_{1}}:=A\cap(B-\boldsymbol{t}_{1})\cap S_{\boldsymbol{t}_{1}}$ is non-null. Clearly, the measure of $A_{\boldsymbol{t}_{1}}$ is at most the measure of $S_{\boldsymbol{t}_{1}}$, which is $\gamma$.
@@
Let $(A_{\boldsymbol{t}})_{\boldsymbol{t}\in T}$ be the sequence returned by the claim. Observe that, for any distinct $\boldsymbol{t},\boldsymbol{t}^{\prime}\in T$, the sets $A_{\boldsymbol{t}}+\boldsymbol{t}$ and $A_{\boldsymbol{t}^{\prime}}+\boldsymbol{t}^{\prime}$ are disjoint; otherwise, by definition of $T$, the sets $A_{\boldsymbol{t}}$ and $A_{\boldsymbol{t}^{\prime}}$ would be at distance at most $20r$ in $G_{d}$, contradicting the fact that $\bigcup_{\boldsymbol{t}\in T}A_{\boldsymbol{t}}$ is $(100r)$-discrete in $G_{d}$. Thus, the sets $A_{\boldsymbol{t}}$ for $\boldsymbol{t}\in T$ form an equidecomposition of the subset
@@ 3.2. Toast Sequences
The following lemma provides a simple construction of a toast sequence which covers $\mathbbm{T}^{k}$ up to a null set. (This will be sufficient for our sketch of the construction of an equidecomposition with Jordan measurable pieces.)
@@ 3.3. Integer-Valued Flow
Let $Z_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,Z_{N}$ be disjoint subsets of a Jordan measurable set $Z\subseteq\mathbbm{T}^{k}$. If, for every real $\varepsilon>0$, there are Jordan measurable subsets $Z_{j}^{\prime}\subseteq Z_{j}$ such that $\lambda(Z\setminus\bigcup_{j=1}^{N}Z_{j}^{\prime})<\varepsilon$, then all of the sets $Z_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,Z_{N}$ are Jordan measurable.
@@ 3.5. Reducing/Analysing the Complexity of the Pieces
Our approach to bounding the upper Minkowski dimension (defined in (1.1)) of the boundaries of the pieces is, for given $\delta>0$, to choose some index $i=i(\delta)$ and run our algorithm to process $J_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,J_{i}$. Although we do not know the final equidecomposition at this moment, we may nonetheless determine for some elements of $A$ to which final piece they will belong. This defines the current partial pieces $A_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,A_{N}$ (that give an equidecomposition of some subset of $A$ to a subset of $B$). Now, we estimate the number of boxes of a regular grid with side length $\delta$ that intersect the currently unassigned part of $A$ or the boundary of any current piece $A_{j}$. (Thus, our dimension estimates depend only on $J_{1},J_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt$ and not on how $K_{1},L_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt$ are built around them.) Recall by Lemma 3.2 that each set $J_{j}$ is a finite union of strips, whose number we can estimate if the sets $X_{i}$ are carefully chosen (namely, as in Lemma 7.6). Also, we can estimate a radius $R$ so that the current pieces $A_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,A_{N}$ can be $R$-locally determined from $A$, $B$ and $J_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,J_{i}$. Thus each $A_{j}$ is a Boolean combination of $(2R+1)^{d}$ translates of each of these sets, and the number of boxes that intersect its boundary $\partial A_{j}$ can be bounded above by the number of boxes intersecting the boundary of at least one of these translates. Furthermore, every other box that may potentially intersect the boundary of a final piece has to be a subset of $A^{\prime}:=A\setminus(\bigcup_{j=1}^{N}A_{j})$ and their number can be upper bounded by $\lambda(A^{\prime})/\delta^{k}$ by the trivial volume argument. Thus we have to control our parameters carefully to get a good balance between the minimal distance between the components of $J_{i}$ (in order to control the measure of the leftover part $A^{\prime}$ of $A$) and their maximum diameter (as our local rule processing $J_{i}$ has to use radius at least as large as this diameter).
@@
In other words, $\phi_{\boldsymbol{u},n\vec{\gamma}}$ sends a unit of flow from $\boldsymbol{u}$ to $(n\vec{\gamma})\cdot\boldsymbol{u}$ along the straight-line path (whose direction is $\vec{\gamma}$ and graph length is $n$). In particular, if $n\vec{\gamma}$ is the zero vector, then the flow $\phi_{\boldsymbol{u},n\vec{\gamma}}$ is identically zero. Also, for an integer $m\geqslant 0$ and a discrete $2^{m+1}$-cube $Q$ containing $\boldsymbol{u}\in{\mathbbm{T}}^{k}$, let
@@
Starting with $f_{0,{\mathcal{G}}_{0}}:=0$ being the zero flow, obtain inductively for $i=1,2,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,m$, the flow $f_{i,{\mathcal{G}}_{i}}$ from $f_{i-1,{\mathcal{G}}_{i-1}}$ by adding
(4.3) $\theta_{i,{\mathcal{G}}_{i}}:=\sum_{Q\in{\mathcal{G}}_{i}}\sum_{\boldsymbol{u}\in Q}\left({\mathbbm{1}}_{A}(\boldsymbol{u})-{\mathbbm{1}}_{B}(\boldsymbol{u})-f_{i-1,{\mathcal{G}}_{i-1}}^{\operatorname{out}}(\boldsymbol{u})\right)\phi_{\boldsymbol{u},Q}.$
Note that, inside each $2^{i}$-cube $Q\in{\mathcal{G}}_{i}$, the increment flow $\theta_{i,{\mathcal{G}}_{i}}$ spreads the current demand error ${\mathbbm{1}}_{A}-{\mathbbm{1}}_{B}-f_{i-1,{\mathcal{G}}_{i-1}}^{\operatorname{out}}$ uniformly inside each congruence class modulo $2^{i-1}$. Of course, if the error ${\mathbbm{1}}_{A}-{\mathbbm{1}}_{B}-f_{i-1,{\mathcal{G}}_{i-1}}^{\operatorname{out}}$ is constant on each cube in $\mathcal{P}(Q)$ for some $Q\in{\mathcal{G}}_{i}$ then $\theta_{i,{\mathcal{G}}_{i}}$ spreads this error evenly inside $Q$. Thus an easy induction on $i=1,2,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,m$ shows that ${\mathbbm{1}}_{A}-{\mathbbm{1}}_{B}-f_{i,{\mathcal{G}}_{i}}^{\operatorname{out}}$ is constant on every cube $Q\in{\mathcal{G}}_{i}$. Moreover, since $f_{i,{\mathcal{G}}_{i}}$ is zero on the edge boundary of $Q\in{\mathcal{G}}_{i}$ (and thus sends no flow out of $Q$), this constant is $\xi(Q)$, that is,
(4.4) ${\mathbbm{1}}_{A}(\boldsymbol{u})-{\mathbbm{1}}_{B}(\boldsymbol{u})-f_{i,{\mathcal{G}}_{i}}^{\operatorname{out}}(\boldsymbol{u})=\xi(Q),\quad\mbox{for all $Q\in{\mathcal{G}}_{i}$ and $\boldsymbol{u}\in Q$}.$
@@
(4.5) $\theta_{i,{\mathcal{G}}_{i}}:=\sum_{Q\in{\mathcal{G}}_{i}}\sum_{C\in\mathcal{P}(Q)}\xi(C)\sum_{\boldsymbol{u}\in C}\phi_{\boldsymbol{u},Q},\quad\mbox{for every $1\leqslant i\leqslant m$}$
@@
It follows from (4.2) and (4.5) that, inside the component $\boldsymbol{v}]_{G_{d}}$, the flow $\theta_{i}$ for each $1\leqslant i\leqslant m$ is the expectation of $\theta_{i,{\mathcal{G}}_{i}}$ over a uniformly random $2^{m}$-grid ${\mathcal{G}}_{m}$, since ${\mathcal{G}}_{i}=\mathcal{P}^{m-i}({\mathcal{G}}_{m})$ is a uniformly random $2^{i}$-grid. Note that, by the linearity of expectation, $f_{m}$ inside $\boldsymbol{v}]_{G_{d}}$ is the expectation of $f_{m,{\mathcal{G}}_{m}}$.
@@ 5.1. Some Preliminaries
Let $D=(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i})$ be a sequence of subsets of $\mathbbm{T}^{k}$, let $b\geqslant 0$ and let $(a_{1},a_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,a_{i-1})$ and $(b_{0},b_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,b_{i-1})$ be sequences of integers such that $b_{0}\geqslant 0$ and
@@
So, suppose that $i\geqslant 2$. If none of the sets $S_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{n}$ is a component of $G_{d}\upharpoonright D_{i-1}$, then the sequence $S_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{n}$ actually certifies that $\boldsymbol{w}\in\mathcal{C}_{b}(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},D_{i})$ and we have that $\operatorname{dist}_{G_{d}}(\boldsymbol{u},\boldsymbol{w})\leqslant b_{i-2}<b_{i-1}$ by induction on $i$.
Next, suppose that there is a unique index $\ell$ such that $S_{\ell}$ is a component of $G_{d}\upharpoonright D_{i-1}$. Let $\boldsymbol{y}$ be a vertex of $S_{\ell}$ at distance at most $2b$ from $S_{\ell-1}$ (or, at distance at most $b$ from $\boldsymbol{u}$ in the case $\ell=1$) and let $\boldsymbol{z}$ be a vertex of $S_{\ell}$ at distance at most $2b$ from $S_{\ell+1}$ (or at distance at most $b$ from $\boldsymbol{w}$ in the case $\ell=n$). The sequences $S_{\ell-1},S_{\ell-2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5ptS_{1}$ and $S_{\ell+1},S_{\ell+2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{n}$ certify that $\boldsymbol{u}$ and $\boldsymbol{w}$, respectively, are in $\mathcal{C}_{b}(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},N_{b}[S_{\ell}])$. So, by the inductive hypothesis and (5.3),
@@
$\leqslant$ $\left(\operatorname{dist}_{G_{d}}(\boldsymbol{u},N_{b}[S_{\ell}])+b\right)+a_{i-1}+\left(\operatorname{dist}_{G_{d}}(N_{b}[S_{\ell}],\boldsymbol{w})+b\right)$
@@
Finally, we consider the remaining case that there are two indices $1\leqslant\ell<\ell^{\prime}\leqslant n$ such that $S_{\ell}$ and $S_{\ell^{\prime}}$ are distinct components of $G_{d}\upharpoonright D_{i-1}$. Choose these indices so that $\ell^{\prime}$-$\ell$ is minimised. Let $\boldsymbol{y}$ be a vertex of $N_{b}[S_{\ell}]$ which is at distance at most $b$ from $S_{\ell+1}$ and let $\boldsymbol{z}$ be a vertex of $N_{b}[S_{\ell^{\prime}}]$ at distance at most $b$ from $S_{\ell^{\prime}-1}$. The sequence $S_{\ell+1},S_{\ell+2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{\ell^{\prime}-1}$ certifies that $\boldsymbol{z}$ is in $\mathcal{C}_{b}(D_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},N_{b}[S_{\ell}])$. By the inductive hypothesis,

### v4

@@ 2.5. Some Auxiliary Structures in $G_{d}$
At certain points during the construction, we have to make arbitrary “tie-breaking” choices in a consistent manner. In doing so, it is useful to endow ${\mathbbm{T}}^{k}$ with a partial order in which each component of $G_{d}$ is totally ordered. We will use the lexicographic ordering where, for distinct $\boldsymbol{u},\boldsymbol{v}\in\mathbbm{T}^{k}$, we write $\boldsymbol{u}\prec_{\operatorname{lex}}\boldsymbol{v}$ to mean $\boldsymbol{u}]_{G_{d}}=[\boldsymbol{v}]_{G_{d}}$ and the first non-zero entry of $\vec{n}$ is positive, where $\vec{n}$ is the (unique) element of ${\mathbbm{Z}}^{d}$ with
@@
Note that the families of subsets of ${\mathbbm{T}}^{k}$ that are Jordan measurable or have boundary of upper Minkowski dimension at most some given constant are both examples of $a$-invariant algebras.
@@
For any $r\geqslant 1$ there exists $\delta>0$ such that every set $X\subseteq\mathbbm{T}^{k}$ whose projection onto the first coordinate has diameter at most $\delta$ is $r$-discrete.
@@
Since $X=C_{1}^{\prime}\cup\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt\cup C_{m}^{\prime}$, it is enough to show that each $C_{i}^{\prime}$ is an r(i1)r(i-1)-local function of $C_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,C_{i}$. We use induction on $i=1,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,m$. This is true for $i=1$ since the set $C_{1}^{\prime}=C_{1}$ is a 0-local function of $C_{1}$. For each $2\leqslant i\leqslant m$, the set $C_{i}^{\prime}$ is a Boolean combination of $C_{i}$ and the sets $C_{1}^{\prime},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,C_{i-1}^{\prime}$ translated by vectors of the form $\sum_{i=1}^{d}n_{i}\boldsymbol{x}_{i}$ where $\|\vec{n}\|_{\infty}\leqslant r$. So, by induction, $C_{i}^{\prime}$ is an r(i1)r(i-1)-local function of $C_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,C_{i}$, as desired.
@@ 2.7. Equidecompositions from Integer-Valued Flows
Start by taking $S_{\boldsymbol{t}_{1}}$ to be a strip of width $\gamma$ such that the intersection
$A_{\boldsymbol{t}_{1}}:=A\cap(B-\boldsymbol{t}_{1})\cap S_{\boldsymbol{t}_{1}}$
is non-null. Clearly, the measure of $A_{\boldsymbol{t}_{1}}$ is at most the measure of $S_{\boldsymbol{t}_{1}}$, which is $\gamma$.
@@
Let $(A_{\boldsymbol{t}})_{\boldsymbol{t}\in T}$ be the sequence returned by the claim. Observe that, for any distinct $\boldsymbol{t},\boldsymbol{t}^{\prime}\in T$, the sets $A_{\boldsymbol{t}}+\boldsymbol{t}$ and $A_{\boldsymbol{t}^{\prime}}+\boldsymbol{t}^{\prime}$ are disjoint; otherwise, by the definition of $T$, the sets $A_{\boldsymbol{t}}$ and $A_{\boldsymbol{t}^{\prime}}$ would be at distance at most $20r$ in $G_{d}$, contradicting the fact that $\bigcup_{\boldsymbol{t}\in T}A_{\boldsymbol{t}}$ is $(100r)$-discrete in $G_{d}$. Thus, the sets $A_{\boldsymbol{t}}$ for $\boldsymbol{t}\in T$ form an equidecomposition of the subset
@@ 3.2. Toast Sequences
The following lemma provides a simple construction of a toast sequence which covers $\mathbbm{T}^{k}$ up to a null set. (This will be sufficient for our construction of an equidecomposition with Jordan measurable pieces.)
@@ 3.3. Integer-Valued Flow
Let $Z_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,Z_{N}$ be pairwise disjoint subsets of a Jordan measurable set $Z\subseteq\mathbbm{T}^{k}$. If, for every real $\varepsilon>0$, there are Jordan measurable subsets $Z_{j}^{\prime}\subseteq Z_{j}$ such that $\lambda(Z\setminus\bigcup_{j=1}^{N}Z_{j}^{\prime})<\varepsilon$, then all of the sets $Z_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,Z_{N}$ are Jordan measurable.
@@ 3.5. Reducing/Analysing the Complexity of the Pieces
Our approach to bounding the upper Minkowski dimension (defined in (1.1)) of the boundaries of the pieces is, for given $\delta>0$, to choose some index $i=i(\delta)$ and run our algorithm to process $J_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,J_{i}$. Although we do not know the final equidecomposition at this moment, we may nonetheless determine for some elements of $A$ to which final piece they will belong. This defines the current partial pieces $A_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,A_{N}$ (that give an equidecomposition of some subset of $A$ to a subset of $B$). Now, we estimate the number of boxes of a regular grid with side length $\delta$ that intersect the currently unassigned part of $A$ or the boundary of any current piece $A_{j}$. (Thus, our dimension estimates depend only on $J_{1},J_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt$ and not on how $K_{1},L_{1},K_{2},L_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt$ are built around them.) Recall by Lemma 3.2 that each set $J_{j}$ is a finite union of strips, whose number we can estimate if the sets $X_{i}$ are carefully chosen (namely, as in Lemma 7.6). Also, we can estimate a radius $R$ so that the current pieces $A_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,A_{N}$ can be $R$-locally determined from $A$, $B$ and $J_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,J_{i}$. Thus each $A_{j}$ is a Boolean combination of $(2R+1)^{d}$ translates of each of these sets, and the number of boxes that intersect its boundary $\partial A_{j}$ can be bounded above by the number of boxes intersecting the boundary of at least one of these translates. Furthermore, every other box that may potentially intersect the boundary of a final piece has to be a subset of $A^{\prime}:=A\setminus(\bigcup_{j=1}^{N}A_{j})$ and their number can be upper bounded by $\lambda(A^{\prime})/\delta^{k}$ by the trivial volume argument. Thus we have to control our parameters carefully to get a good balance between the minimal distance between the components of $J_{i}$ (in order to control the measure of the leftover part $A^{\prime}$ of $A$) and their maximum diameter (as our local rule processing $J_{i}$ has to use radius at least as large as this diameter).
@@
In other words, $\phi_{\boldsymbol{u},n\vec{\gamma}}$ sends a unit of flow from $\boldsymbol{u}$ to $(n\vec{\gamma})\cdot\boldsymbol{u}$ along the straight-line path (whose direction is $\vec{\gamma}$ and graph length is $|n|$). In particular, if $n\vec{\gamma}$ is the zero vector, then the flow $\phi_{\boldsymbol{u},n\vec{\gamma}}$ is identically zero. Also, for an integer $m\geqslant 0$ and a discrete $2^{m+1}$-cube $Q$ containing $\boldsymbol{u}\in{\mathbbm{T}}^{k}$, let
@@
Starting with $f_{{\mathcal{G}}_{0}}:=0$ being the zero flow, obtain inductively for $i=1,2,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,m$, the flow $f_{{\mathcal{G}}_{i}}$ from $f_{{\mathcal{G}}_{i-1}}$ by adding
(4.3) $\theta_{{\mathcal{G}}_{i}}:=\sum_{Q\in{\mathcal{G}}_{i}}\sum_{\boldsymbol{u}\in Q}\left({\mathbbm{1}}_{A}(\boldsymbol{u})-{\mathbbm{1}}_{B}(\boldsymbol{u})-f_{{\mathcal{G}}_{i-1}}^{\operatorname{out}}(\boldsymbol{u})\right)\phi_{\boldsymbol{u},Q}.$
Note that $f_{{\mathcal{G}}_{i}}$ and $\theta_{{\mathcal{G}}_{i}}$ do not depend on $m$ (for $m\geqslant i$). Inside each $2^{i}$-cube $Q\in{\mathcal{G}}_{i}$, the increment flow $\theta_{{\mathcal{G}}_{i}}$ spreads the current demand error ${\mathbbm{1}}_{A}-{\mathbbm{1}}_{B}-f_{{\mathcal{G}}_{i-1}}^{\operatorname{out}}$ uniformly inside each congruence class modulo $2^{i-1}$. Of course, if the error ${\mathbbm{1}}_{A}-{\mathbbm{1}}_{B}-f_{{\mathcal{G}}_{i-1}}^{\operatorname{out}}$ is constant on each cube in $\mathcal{P}(Q)$ for some $Q\in{\mathcal{G}}_{i}$ then $\theta_{{\mathcal{G}}_{i}}$ spreads this error evenly inside $Q$. Thus an easy induction on $i=1,2,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,m$ shows that ${\mathbbm{1}}_{A}-{\mathbbm{1}}_{B}-f_{{\mathcal{G}}_{i}}^{\operatorname{out}}$ is constant on every cube $Q\in{\mathcal{G}}_{i}$. Moreover, since $f_{{\mathcal{G}}_{i}}$ is zero on the edge boundary of $Q\in{\mathcal{G}}_{i}$ (and thus sends no flow out of $Q$), this constant is $\xi(Q)$, that is,
(4.4) ${\mathbbm{1}}_{A}(\boldsymbol{u})-{\mathbbm{1}}_{B}(\boldsymbol{u})-f_{{\mathcal{G}}_{i}}^{\operatorname{out}}(\boldsymbol{u})=\xi(Q),\quad\mbox{for all $Q\in{\mathcal{G}}_{i}$ and $\boldsymbol{u}\in Q$}.$
@@
(4.5) $\theta_{{\mathcal{G}}_{i}}:=\sum_{Q\in{\mathcal{G}}_{i}}\sum_{C\in\mathcal{P}(Q)}\xi(C)\sum_{\boldsymbol{u}\in C}\phi_{\boldsymbol{u},Q},\quad\mbox{for every $1\leqslant i\leqslant m$}$
@@
It follows from (4.2) and (4.5) that, inside the component $\boldsymbol{v}]_{G_{d}}$, the flow $\theta_{i}$ for each $1\leqslant i\leqslant m$ is the expectation of $\theta_{{\mathcal{G}}_{i}}$ over a uniformly random $2^{m}$-grid ${\mathcal{G}}_{m}$, since ${\mathcal{G}}_{i}=\mathcal{P}^{m-i}({\mathcal{G}}_{m})$ is a uniformly random $2^{i}$-grid. Note that, by the linearity of expectation, $f_{m}$ inside $\boldsymbol{v}]_{G_{d}}$ is the expectation of $f_{{\mathcal{G}}_{m}}$.
@@ 5.1. Some Preliminaries
Let $D=(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i})$ be a sequence of subsets of $\mathbbm{T}^{k}$, let $b\geqslant 0$, and let $(a_{1},a_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,a_{i-1})$ and $(b_{0},b_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,b_{i-1})$ be sequences of non-negative integers such that
@@
So, suppose that $i\geqslant 2$. If none of the sets $S_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{n}$ is a component of $G_{d}\upharpoonright D_{i-1}$, then the sequence $S_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{n}$ actually certifies that $\boldsymbol{w}\in\mathcal{C}_{b}(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},\boldsymbol{u})$ and we have that $\operatorname{dist}_{G_{d}}(\boldsymbol{u},\boldsymbol{w})\leqslant b_{i-2}\leqslant b_{i-1}$ by induction on $i$.
Next, suppose that there is a unique index $\ell$ such that $S_{\ell}$ is a component of $G_{d}\upharpoonright D_{i-1}$. Let $\boldsymbol{y}$ be a vertex of $S_{\ell}$ at distance at most $2b$ from $S_{\ell-1}$ (or, at distance at most $b$ from $\boldsymbol{u}$ in the case $\ell=1$) and let $\boldsymbol{z}$ be a vertex of $S_{\ell}$ at distance at most $2b$ from $S_{\ell+1}$ (or at distance at most $b$ from $\boldsymbol{w}$ in the case $\ell=n$). The sequences $S_{\ell-1},S_{\ell-2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5ptS_{1}$ and $S_{\ell+1},S_{\ell+2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{n}$ certify respectively that $\boldsymbol{u}\in\mathcal{C}_{b}(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},N_{b}[\boldsymbol{y}])$ and $\boldsymbol{w}\in\mathcal{C}_{b}(D_{1},D_{2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},N_{b}[\boldsymbol{z}])$. So, by the inductive hypothesis and (5.3),
@@
$\leqslant$ $\left(\operatorname{dist}_{G_{d}}(\boldsymbol{u},N_{b}[\boldsymbol{y}])+b\right)+a_{i-1}+\left(\operatorname{dist}_{G_{d}}(N_{b}[\boldsymbol{z}],\boldsymbol{w})+b\right)$
@@
Finally, we consider the remaining case that there are two indices $1\leqslant\ell<\ell^{\prime}\leqslant n$ such that $S_{\ell}$ and $S_{\ell^{\prime}}$ are distinct components of $G_{d}\upharpoonright D_{i-1}$. Choose these indices so that $\ell^{\prime}$-$\ell$ is minimised. Let $\boldsymbol{y}$ be a vertex of $N_{b}[S_{\ell}]$ which is at distance at most $b$ from $S_{\ell+1}$ and let $\boldsymbol{z}$ be a vertex of $N_{b}[S_{\ell^{\prime}}]$ at distance at most $b$ from $S_{\ell^{\prime}-1}$. The sequence $S_{\ell+1},S_{\ell+2},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,S_{\ell^{\prime}-1}$ certifies that $\boldsymbol{z}$ is in $\mathcal{C}_{b}(D_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,D_{i-2},\{\boldsymbol{y}\})$. By the inductive hypothesis,